Bastian Gruber

Professional LinkedIn: linkedin.com/in/bastiangruber Project: rustwebdevelopment.com Work GitHub: github.com/gruberb SourceHut: git.sr.ht/ gruberb

SUMMARY

Passionate Software Engineer and published Author, worked in many roles (Tech Lead, Customer Facing, Systems Engineer, Frontend and Backend) in multiple companies (StartUps with 3-8 people, small businesses with around 50 people and large scale enterprise from 200 to 5.000 people). I have 9 years experience after my Computer Science studies, and 5 years before where I founded my own web shop as acting CTO. I am used to create MVPs as well as large-scale, high-scalability web services. I am considered a great team player, and lover of software.

PROFESSIONAL EXPERIENCE

Sr. Solutions Architect (Professional Services) - Full-time June 2020 - now Twilio Inc., Remote Twilio just started to create their Professional Services team in EMEA and I was the second SA joining their team. The goal is to support and enable small to large scale customers using the Twilio APIs, help them integrate the services, do code reviews and review their architecture.

• Acted as the Technical Lead

• Involved in hiring and onboarding

• Creating architectural specifications

• Develop reusable deliverables for customers

• Leading customers through a technical solution

• Code review

• Setting up sales and feedback cycles

Author/Blogger - Contract Several Tech Publications and Blogs January 2008 - now

• Publishing the book ”Rust Web Development” with Manning

• Writing articles for the german Macworld, having my own OS X/macOS column

• Writing tech articles for the largest publisher (Heise) in Germany (iX, mac&I, c’t)

• Writing for blogs like mobilegeeks.de and stadt-bremerhaven.de

• Publishing a book for bookboon.com ”OS X for beginners”

System Engineer (Rust) - Contract September 2019 - May 2020 Kraken Digital Currency Exchange, Remote (San Francisco) Part of the first Rust team of the company, replacing the core trading engine with Rust. Worked alongside 12 highly experienced Software Developers from all over the world. Refactor code from PHP to Rust, wrote new features in Rust in a well-tested highly secure manner.

• Re-wrote several PHP services in Rust

• Built new features in Rust directly

• Developed programs to sustain several thousand requests a second

Software Engineer - Contract September 2018 – February 2019 OSRAM, Berlin (Germany) Setting up the authentication part of a large scale IoT platform. Using Go and NodeJS to develop ID management, offering APIs to in-house and third-party developers. We had several junior developers in the team where my job was to mentor and guide them through decision making.

• Set up the authentication part of a large-scale IoT platform

• Analysed problem areas and found technical solutions, either in-house or from the third parties

• Enabled clients authenticate with an IoT system and access the platform in a secure and fast way

• Developed ID management with Go and NodeJS, offering APIs to in-house and third-party developers

• Mentored and guided junior developers

Software Engineer - Contract June 2018 – August 2018 Sevenval Technologies GmbH, Berlin (Germany) The company hired me to set up a better structure in their day-to-day development process, clean up their code base and make architectural decisions for their current product they were developing. I acted as a Tech Lead, enabled Scrum processes, wrote new code in NodeJS (TypeScript) and helped improve the frontend code as well (Angular 6).

Software Engineer - Contract February 2018 – May 2018 Körber Digital, Berlin (Germany) I joined as one of the first team members, helping setting up the tech stack for the frontend and building microserives for the Backend. The main focus was building infrastructure for an IoT product with close to real-time data. Reading time series sensor data from machines and distribute it to different clients. The job was to set up an infrastructure which can support real-time data and creating an easy to use RESTful API.

Travel and Code - Personal August 2017 – January 2018 North America Traveled through Canada and the United States. Visited MeetUps in San Francisco, Portland, Toronto and Montreal. Did Blockchain workshops and discovered tech scenes. I camped along the way, hiked and biked around nature. This time off was meant to reflect on life and pause for a moment.

Software Engineer - Contract April 2017 – July 2017 Ergo Group AG, Berlin (Germany) Developing in-house tools for managing clients and routing for appoinments. Also creating new customer facing products for managing insurance contracts online. I joined as one of the first team members in the Ergo Digital Hub and helped forming processes in addition to guiding team members and mentoring junior delvopers. The work was done in JavaScript (React and NodeJS).

Software Engineer - Contract November 2016 – March 2017 GIATA GmbH, Berlin (Germany) GIATA has a huge amount of travel data from hotels to other POIs. The previous infrastructure was written in PHP. The goals was to create a new environment consisting of micro services deployed via Docker. I was part of a 4 person team who taught the existing developers Scrum, JavaScript and set up the new backend and frontend.

Software Engineer - Contract October 2016 – December 2016 PaperHive, Berlin (Germany) PaperHive is a platform to discuss scientific papers and share them. Newly created papers can also be shared in private groups prior publishing. The goal was to help the team setting up private messaging channels and refactoring the existing NodeJS backend.

Senior Frontend Developer - Full-time March 2016 – September 2016 glispa GmbH, Berlin (Germany) I was hired to develop the in-house application further and refactor parts of the old code base. I worked together with 6 frontend developers to create a dynamic dashboard for in-house anlytics but also customer facing frontends to buy ad-banners for their products.

Software Engineer - Full-time August 2015 – Februar 2016 smartB GmbH, Berlin (Germany) I worked in a small and very agile team to create an application which lets business owner see their energy consumption. We worked heavily with pair programming and a developer was involved in the whole stack (JavaScript, Python).

Software Engineer - Contract March 2015 – July 2015 blogfoster, Berlin (Germany) I joined as the first developers to re-write an existing PHP application in the backend and a Backbone frontend. We were using NodeJS and React to re-build the application (a customer-facing dashboard to buy banners for blogs).

Frontend Developer - Full-time December 2013 – March 2015 Dalia Research, Berlin (Germany) Dalia Research was founded in 2013 in Berlin. We are working with a small group of 8 people on products which change the way delivering questions to people around the globe. As the only Frontend Developer I was responsible for responsive web interfaces across all browsers and devices. I joined as the first Frontend Developer and was responsible for setting up the tech stack, build pipelines and logging systems.

Scholarship Student - Full-time September 2010 – February 2014 Allianz Deutschland AG/ AMOS SE, Munich (Germany) This scholarship was an intense 4 years program where I was working in my semester breaks and also parallel during the semester at TH Ingolstadt. I was part of different teams (large-scale batch processing on the mainframe, Java client software, web).

Co-Founder/CTO - Full-time January 2008 – November 2010 Sunny Dust GbR, Pappenheim (Germany) Founded a company to help companies creating their web presence. We created websites in Flash and HTML, and also created and printed business cards and flyers. Working closely with the customers at this young age helped me to understand the business site and the processes which are involved to create a product and actually selling it.

EDUCATION

Bachelor of Science, Computer Science TH Ingolstadt February 2014

PUBLICATIONS

• ”Practical Web Development with Rust” - Manning Publications

• ”OS X Lion for beginners” - bookboon.com

VOLUNTEER EXPERIENCE

Mentor - Geek Girls Carrot Hackathon Founder - ”Writers Lounge Berlin” MeetUp Founder/Teacher - ”Web Development for Artists” MeetUp Founder - ”Rust and Tell Berlin” MeetUp

INTERESTS Piano, Running, Philosophy